OpenStudy (richyw):

Superposition of states. How can show that these wavefunctions do or do not satisfy the time-independant Schrodinger equation?

OpenStudy (richyw):

If I have two normalized eigenkets of the hamiltonian. \(\left| a\right\rangle\) and \(\left| b\right\rangle\), with eigenvalues \(E_a\) and \(E_b\), how would I show that its corresponding wavefunction satisfies the shrodinger equation? what about the wavefunction for a linear combination of the two states?

OpenStudy (kainui):

Do you have the time-independent Schrodinger equation there? If so, you essentially just plug it in and find out if you get 1=1.

OpenStudy (richyw):

plug what in?

OpenStudy (anonymous):

\[H|a>=E_a|a>\] \[H|b>=E_b|b>\] \[H(|a>+|b>)=E_a|a>+E_b|b>\neq (E_a+E_b)(|a>+|b>)\] then \[H(\alpha|a>+\beta|b>)=\alpha E_a|a>+\beta E_b|b>\] for some alpha and beta, such that \[\alpha=1+E_b/E_a\] \[\beta=1+E_a/E_b\] then you can have \[H(\alpha)|a>+\beta|b>=(E_a+E_b)(|a>+|b>)\] I am not sure !